ADNAN MUHAMMED

Grade 12,

Let Sn denote the sum of first n terms of an A.P.If S4/S2=4,then Sn+1/Sn is equal to-----------------

Let Sn denote the sum of first n terms of an A.P.If S4/S2=4,then Sn+1/Sn is equal to-----------------


 

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sri

let first term is a and common difference is d

 S4 = 2(2a + 3d)

 S2 = 2a +d

 given 2(2a + 3d)/(2a +d) = 4

          2a + 3d = 4a + 2d

             d = 2a ...........1

now Sn+1 /Sn  = (n+1)/2 [2a+ nd] / n/2[2a +(n-1)d]

                      = (n+1)[ 1+n]/n[n]

                      =(1+1/n)2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin